# Self interacting scalar field theory in general curved spacetimes at zero and finite temperature revisited

### v1

@@ 6 Two loop effective potential in curved spacetime
We now wish to extend some of our earlier results to two loop order. The effective action at this order reads ($\hbar=1$)
$-\frac{\lambda}{8}\int\sqrt{-g}d^{d}xG^{2}(x,x)+\frac{1}{12}\int\sqrt{-g}\sqrt{-g^{\prime}}d^{d}xd^{d}x^{\prime}(\eta+\lambda\phi)^{2}G^{3}(x,x^{\prime})+\ {\rm counterterm~{}contributions}$ (53)
Note that the second or the sunset integral is quadratic in the couplings. Thus we may expect a priori that the first, i.e. the double bubble integral to make the leading contributions at two loop, also supported by our earlier argument of 39. The corresponding 1PI vacuum diagrams can be seen 14.
@@
Let us now compute the diagrams of 14. In order to keep our expressions simple, we shall also make a linear in curvature approximation in the following. We evaluate the zero temperature effective potential first, show the renormalisation and then turn on the temperature. We will see that the latter is divergence free, as expected. Finally we shall present the finite part of the ${\cal O}(\lambda)$ double bubble diagram at finite temperature up to quadratic order in the curvature.
@@
The double bubble diagram is given by the first of 14. We have at $d=4-\epsilon$,
@@
$\left.+2\left(\ln{\frac{m^{2}_{1}}{4\pi\mu^{2}}}\right)^{2}+\psi(1)\psi(2)+\frac{1}{2}\left(\frac{\pi^{2}}{3}+\psi^{2}(2)-\psi^{\prime}(2)\right)+\frac{1}{2}\left(\frac{\pi^{2}}{3}+\psi^{2}(1)-\psi^{\prime}(1)\right)+2\ln{\frac{4\pi\mu^{2}}{m^{2}_{1}}}\left(\psi(1)+\psi(2)\right)\right)$
@@
$\psi^{\prime}(x)=\frac{\Gamma^{\prime\prime}(x)}{\Gamma(x)}-\frac{(\Gamma^{\prime}(x))^{2}}{\Gamma^{2}(x)}$
@@
The sunset integral, up to linear in $R$ terms reads
$-\frac{\left(\eta+\lambda\phi\right)^{2}m^{2}_{1}}{2(4\pi)^{4}}\left(\frac{\mu^{-2\epsilon}}{\epsilon^{2}}+\frac{\mu^{-\epsilon}}{\epsilon}\left(\frac{3}{2}+\psi(1)-\ln{\frac{m^{2}_{1}}{4\pi\mu^{2}}}\right)+2\left(\ln\frac{4\pi\mu^{2}}{m^{2}_{1}}\right)^{2}+2\left(\psi^{\prime}(1)+\psi^{2}(1)\right)\right.$
$\left.+4\psi(1)\ln\frac{4\pi\mu^{2}}{m^{2}_{1}}+\ln\frac{4\pi\mu^{2}}{m^{2}_{1}}+\psi(1)-\frac{1}{2}\right)$
$+\frac{\left(\eta+\lambda\phi\right)^{2}}{2}\frac{\left(\frac{1}{6}-\xi\right)R}{(4\pi)^{4}}\left(\frac{\mu^{-2\epsilon}}{\epsilon^{2}}+\frac{\mu^{-\epsilon}}{\epsilon}\left(\frac{1}{2}+\psi(1)-\ln{\frac{m^{2}_{1}}{4\pi\mu^{2}}}\right)+2\left(\ln\frac{4\pi\mu^{2}}{m^{2}_{1}}\right)^{2}+2\left(\psi^{\prime}(1)+\psi^{2}(1)\right)\right.$
$\left.+4\psi(1)\ln\frac{4\pi\mu^{2}}{m^{2}_{1}}+\ln\frac{4\pi\mu^{2}}{m^{2}_{1}}+\psi(1)-\frac{1}{2}\right)$ (55)
@@
$\frac{1}{2}\left(\delta m^{2}+\frac{1}{2}\delta\lambda\phi^{2}+\delta\eta\phi+\delta\xi R\right)\times\int\frac{d^{d}k}{(2\pi)^{d}}\left(\frac{1}{k^{2}+m^{2}_{1}}+\frac{\left(\frac{1}{6}-\xi\right)R}{(k^{2}+m^{2}_{1})^{2}}\right)$ (56)
where $\delta m^{2}$, $\delta\lambda$ and $\delta\xi$ can be seen in 18. We next add the divergent terms coming from 54, 55, 56, in order to find out the divergence of the effective potential
@@
$\left.+\eta\lambda^{2}\phi^{3}\left(2\psi(2)-2\psi(1)-3\right)+\frac{\eta^{2}}{2}\left(\frac{1}{6}-\xi\right)R+\eta\lambda\phi\left(\frac{1}{6}-\xi\right)R+\frac{\lambda^{2}\phi^{2}}{2}\left(\frac{1}{6}-\xi\right)R\right)$ (57)
@@
$\delta\Lambda=-\frac{\mu^{-2\epsilon}\left(m^{4}\lambda+2m^{2}\eta^{2}\right)}{2(4\pi)^{4}\epsilon^{2}}-\frac{2\mu^{-\epsilon}m^{2}\eta^{2}\left(\psi(2)-\psi(1)-\frac{3}{2}\right)}{2(4\pi)^{4}\epsilon}$
$\delta\kappa=\frac{\mu^{-2\epsilon}\left(2m^{2}\lambda+\eta^{2}\right)}{2(4\pi)^{4}\epsilon^{2}}\left(\frac{1}{6}-\xi\right)-\frac{\eta^{2}\mu^{-\epsilon}}{4(4\pi)^{4}\epsilon}\left(\frac{1}{6}-\xi\right)$
$\delta m^{2}=\frac{\left(4m^{2}\lambda^{2}+\frac{7}{2}\lambda\eta^{2}\right)\mu^{-2\epsilon}}{(4\pi)^{4}\epsilon^{2}}+\frac{2m^{2}\lambda^{2}\mu^{-\epsilon}}{(4\pi)^{4}\epsilon}\left(\frac{3}{4}\psi(2)-\frac{1}{2}\psi(1)-\frac{3}{4}\right)+\frac{\eta^{2}\lambda\mu^{-\epsilon}}{(4\pi)^{4}\epsilon}\left(\frac{5}{2}\psi(2)-\frac{5}{2}\psi(1)-\frac{15}{4}\right)$
$\delta\xi=-\frac{\mu^{-2\epsilon}\lambda^{2}}{(4\pi)^{4}\epsilon^{2}}\left(\frac{1}{6}-\xi\right)+\frac{\lambda^{2}\mu^{-\epsilon}}{2(4\pi)^{4}\epsilon}\left(\frac{1}{6}-\xi\right)$
$\delta\tau=\frac{\left(4m^{2}\eta\lambda+\eta^{3}\right)\mu^{-2\epsilon}}{2(4\pi)^{4}\epsilon^{2}}+\frac{m^{2}\eta\lambda\mu^{-\epsilon}}{2(4\pi)^{4}\epsilon}\left(2\psi(2)-2\psi(1)-3\right)+\frac{\eta^{3}\mu^{-\epsilon}}{2(4\pi)^{4}\epsilon}\left(\psi(2)-\psi(1)-\frac{3}{2}\right)$
$\delta\gamma=-\frac{4\eta\lambda\mu^{-2\epsilon}}{2(4\pi)^{4}\epsilon^{2}}\left(\frac{1}{6}-\xi\right)-\frac{\eta\lambda\mu^{-\epsilon}}{2(4\pi)^{4}\epsilon}\left(\frac{1}{6}-\xi\right)$
$\delta\eta=\frac{9\mu^{-2\epsilon}\eta\lambda^{2}}{(4\pi)^{4}\epsilon^{2}}+\frac{3\mu^{-\epsilon}\eta\lambda^{2}}{(4\pi)^{4}\epsilon}\left(2\psi(2)-2\psi(1)-3\right)$
$\delta\lambda=\frac{9\lambda^{3}\mu^{-2\epsilon}}{(4\pi)^{4}\epsilon^{2}}+\frac{6\lambda^{3}\mu^{-\epsilon}}{(4\pi)^{4}\epsilon}\left(\psi(2)-\psi(1)-\frac{3}{2}\right)$ (58)
Apart from some multiplicative numerical factors, the above result agrees with that of [4].
@@
$\left.+\ln\frac{4\pi\mu^{2}}{m^{2}_{1}}+\psi(1)-\frac{1}{2}\right)$ (59)
@@
$\left.+2\ln{\frac{4\pi\mu^{2}}{m^{2}_{1}}}\left(\psi(1)+\psi(2)\right)\right)+{\cal O}(\lambda^{2})$ (60)
where the one loop effective potential is given by 20. However, we note that the one loop effective potential contains terms quadratic in the curvature as well, whereas in the two loop computation, we have retained terms only linear in the curvature. Hence we shall add with 60 the second order in curvature contribution coming from the ${\cal O}(\lambda)$ double bubble, reading
$V_{{\rm eff},{\cal O}(R^{2})}^{2-{\rm loop}}=-\frac{\lambda}{8(4\pi)^{4}}\left({2f_{1}}\left(\ln{\frac{m^{2}_{1}}{4\pi\mu^{2}}-\psi(2)}\right)+\left(\frac{1}{6}-\xi\right)^{2}R^{2}\left(4\psi(1)\ln{\frac{4\pi\mu^{2}}{m_{1}^{2}}}+\frac{3}{2}\left(\ln{\frac{4\pi\mu^{2}}{m_{1}^{2}}}\right)^{2}+\frac{\pi^{2}}{3}+2\psi^{2}(1)-\psi^{\prime}(1)\right)\right)$ (61)
We shall not go into the detail of the derivation for the above. The above contribution only required some gravitational counterterms for the sake of renormalisation. We have plotted 60 plus 61 in 15 for the de Sitter spacetime. Note that compared to the zero temperature one loop result, 43, there is no qualitatively new change. In particular for $\bar{\eta}=0$, the symmetry breaking feature remains intact.
@@
Figure 15: The two loop double bubble effective potential, 60, at vanishing temperature in the de Sitter spacetime for values $\bar{\eta}=0.447,\ \lambda=0.1,\ \xi=0.125,\ \Lambda=10^{-52}{\rm m}^{-2}$. Note that for the quartic self interaction, the symmetry breaking feature remains intact compared to that of one loop, 5.
@@ 6.1 Two loop at finite temperature
$\frac{1}{k^{2}+m^{2}}\rightarrow\frac{1}{k^{2}+m^{2}}+2\pi n_{B}({|k^{0}|})\delta(k^{2}-m^{2})=\frac{1}{k^{2}+m^{2}}+n_{1}(\beta,k)\qquad({\rm say})$ (62)
And,
$\frac{1}{\left(k^{2}+m^{2}\right)^{2}}\rightarrow\frac{1}{\left(k^{2}+m^{2}\right)^{2}}+n_{2}(\beta,k)$ (63)
@@
$\int\frac{d^{d}k}{(2\pi)^{d}}\frac{1}{k^{2}+m^{2}}\rightarrow\int\frac{d^{d}k}{(2\pi)^{d}}\frac{1}{k^{2}+m^{2}}+S_{1}(\beta)=-m^{2}\left(\frac{2\mu^{-\epsilon}}{\epsilon}+\psi(2)+\ln{\frac{4\pi\mu^{2}}{m^{2}}}\right)+S_{1}(\beta)$ (64)
@@
$\int\frac{d^{d}k}{(2\pi)^{d}}\frac{1}{(k^{2}+m^{2})^{2}}\rightarrow\int\frac{d^{d}k}{(2\pi)^{d}}\frac{1}{(k^{2}+m^{2})^{2}}+S_{2}(\beta)=\left(\frac{2\mu^{-\epsilon}}{\epsilon}+\psi(1)+\ln{\frac{4\pi\mu^{2}}{m^{2}}}\right)+S_{2}(\beta)$ (65)
@@
$\frac{\left(\eta+\lambda\phi\right)^{2}}{12}\left(\int\frac{d^{d}k_{1}}{(2\pi)^{d}}\frac{d^{d}k_{2}}{(2\pi)^{d}}\left(\frac{1}{k_{1}^{2}+m^{2}_{1}}+n_{1}(k_{1})\right)\left(\frac{1}{k_{2}^{2}+m^{2}_{1}}+n_{1}(\beta,k_{2})\right)\left(\frac{1}{(k_{1}+k_{2})^{2}+m^{2}_{1}}+n_{1}(\beta,k_{1}+k_{2})\right)\right.$
$\left.+\left(\frac{1}{6}-\xi\right)R\int\frac{d^{d}k_{1}}{(2\pi)^{d}}\frac{d^{d}k_{2}}{(2\pi)^{d}}\left(\frac{1}{k_{1}^{2}+m^{2}_{1}}+n_{1}(\beta,k_{1})\right)\left(\frac{1}{(k_{2}^{2}+m^{2}_{1})^{2}}+n_{2}(k_{2})\right)\left(\frac{1}{(k_{1}+k_{2})^{2}+m^{2}_{1}}+n_{1}(\beta,k_{1}+k_{2})\right)\right.$
@@
$+\frac{\left(\eta+\lambda\phi\right)^{2}}{12}\left(\frac{6\mu^{-\epsilon}}{\epsilon}S_{1}(\beta)+\left(\frac{1}{6}-\xi\right)R\frac{6\mu^{-\epsilon}}{\epsilon}S_{2}(\beta)\right)+\rm{Finite\,\,terms}$ (67)
@@
$+\frac{\mu^{-\epsilon}}{2(4\pi)^{2}\epsilon}\left(-m^{2}\lambda-\eta^{2}-\frac{3}{2}\lambda^{2}\phi^{2}-3\eta\lambda+\left(\frac{1}{6}-\xi\right)\lambda R\right)\left(S_{1}(\beta)+\left(\frac{1}{6}-\xi\right)RS_{2}(\beta)\right)$ (68)
We now collect and add all the divergent terms from 66, 67 and 68. We obtain after some calculations,
@@
$\left.+\left(\frac{1}{6}-\xi\right)R\int\frac{d^{d}k_{1}}{(2\pi)^{d}}\frac{d^{d}k_{2}}{(2\pi)^{d}}\frac{1}{k_{1}^{2}+m^{2}_{1}}\frac{1}{k_{2}^{2}+m^{2}_{1}}\frac{1}{((k_{1}+k_{2})^{2}+m^{2}_{1})^{2}}\right)$ (69)
@@
$\left.-\frac{15m_{1}^{6}\beta^{4}\zeta(7)}{3072\pi^{8}}+\frac{15m_{1}^{7}\beta^{5}\zeta(7)}{24576\pi^{9}}-\frac{15m_{1}^{7}\beta^{5}\zeta(7)}{393216\pi^{9}}+\frac{3m_{1}^{8}\beta^{6}\zeta(3)\zeta(5)}{524288\pi^{10}}+\frac{m_{1}^{8}\beta^{6}\zeta(3)\zeta(5)}{262144\pi^{10}}\right)+{\cal O}(R^{2})$ (70)
where $V^{(2)}_{\rm{db}}$ is the corresponding zero temperature part. Inclusion of ${\cal O}(R^{2})$ terms to the above result is easy, but we shall not go into such computation here.

### v2

@@ 6 Two loop effective potential in curved spacetime
We now wish to extend some of our earlier results to two loop order. As we have stated towards the end of 5, the following calculations are supposed to be a bridge between the one loop results and the non-perturbative computations we wish to do in future. Also, we wish to verify and demonstrate that the field renormalisation counterterms are indeed independent of the spactime curvature. The two loop renormalisation of $\phi^{4}$-theory in general curved spacetimes was first done in [37]. We note that at two loop, there can be non-local contributions from the integrals of the sunset diagram, 15(b). This originates from the integration of the non-local part of the self energy over the entire spacetime. However, as we have discussed, since the current formalism we are using is essentially based upon local Lorentz invariance and hence suitable at small scales only, any such non-local results could be misleading. Therefore we wish to focus only on the local contributions from the vacuum graphs.
The effective action at this order reads ($\hbar=1$)
$-\frac{\lambda}{8}\int\sqrt{-g}d^{d}xG^{2}(x,x)+\frac{1}{12}\int\sqrt{-g}\sqrt{-g^{\prime}}d^{d}xd^{d}x^{\prime}(\eta+\lambda\phi)^{2}G^{3}(x,x^{\prime})+\ {\rm counterterm~{}contributions}.$ (56)
Note that the second or the sunset integral is quadratic in the couplings. Thus we may expect a priori that the first, i.e. the double bubble integral to make the leading contributions at two loop, also supported by our earlier argument of 40. Note also that the double bubble yields a purely local contribution. The corresponding 1PI vacuum diagrams can be seen 15.
@@
Let us now compute the diagrams of 15. In order to keep our expressions simple, we shall also make a linear in curvature approximation in the following. We evaluate the zero temperature effective potential first, show the renormalisation and then turn on the temperature. We will see that the latter is divergence free, as expected. Finally, we shall present the finite part of the ${\cal O}(\lambda)$ double bubble diagram at finite temperature up to quadratic order in the curvature.
@@
The double bubble diagram is given by the first of 15. We have at $d=4-\epsilon$,
@@
$\left.+2\left(\ln{\frac{m^{2}_{1}}{4\pi\mu^{2}}}\right)^{2}+\psi(1)\psi(2)+\frac{1}{2}\left(\frac{\pi^{2}}{3}+\psi^{2}(2)-\psi^{\prime}(2)\right)+\frac{1}{2}\left(\frac{\pi^{2}}{3}+\psi^{2}(1)-\psi^{\prime}(1)\right)+2\ln{\frac{4\pi\mu^{2}}{m^{2}_{1}}}\left(\psi(1)+\psi(2)\right)\right),$
@@
$\psi^{\prime}(x)=\frac{\Gamma^{\prime\prime}(x)}{\Gamma(x)}-\frac{(\Gamma^{\prime}(x))^{2}}{\Gamma^{2}(x)}.$
@@
As we have mentioned at the beginning of this section, for the sunset diagram we will consider only its local part, as any non-local part could yield misleading results in this local or UV effective formalism. The corresponding integral, up to linear in curvature terms reads
$\frac{\left(\eta+\lambda\phi\right)^{2}}{12}\left(\int\frac{d^{d}k_{1}}{(2\pi)^{d}}\frac{d^{d}k_{2}}{(2\pi)^{d}}\frac{1}{k_{1}^{2}+m^{2}_{1}}\frac{1}{k_{2}^{2}+m^{2}_{1}}\frac{1}{(k_{1}+k_{2})^{2}+m^{2}_{1}}\right.$
$\left.+\left(\frac{1}{6}-\xi\right)R\int\frac{d^{d}k_{1}}{(2\pi)^{d}}\frac{d^{d}k_{2}}{(2\pi)^{d}}\frac{1}{k_{1}^{2}+m^{2}_{1}}\frac{1}{k_{2}^{2}+m^{2}_{1}}\frac{1}{((k_{1}+k_{2})^{2}+m^{2}_{1})^{2}}\right)=-\frac{\left(\eta+\lambda\phi\right)^{2}m^{2}_{1}}{2(4\pi)^{4}}\left(\frac{\mu^{-2\epsilon}}{\epsilon^{2}}+\frac{\mu^{-\epsilon}}{\epsilon}\left(\frac{3}{2}+\psi(1)-\ln{\frac{m^{2}_{1}}{4\pi\mu^{2}}}\right)\right.$
$\left.+2\left(\ln\frac{4\pi\mu^{2}}{m^{2}_{1}}\right)^{2}+2\left(\psi^{\prime}(1)+\psi^{2}(1)\right)+4\psi(1)\ln\frac{4\pi\mu^{2}}{m^{2}_{1}}+\ln\frac{4\pi\mu^{2}}{m^{2}_{1}}+\psi(1)-\frac{1}{2}\right)+\frac{\left(\eta+\lambda\phi\right)^{2}}{2}\frac{\left(\frac{1}{6}-\xi\right)R}{(4\pi)^{4}}\left(\frac{\mu^{-2\epsilon}}{\epsilon^{2}}\right.$
$\left.+\frac{\mu^{-\epsilon}}{\epsilon}\left(\frac{1}{2}+\psi(1)-\ln{\frac{m^{2}_{1}}{4\pi\mu^{2}}}\right)+2\left(\ln\frac{4\pi\mu^{2}}{m^{2}_{1}}\right)^{2}+2\left(\psi^{\prime}(1)+\psi^{2}(1)\right)+4\psi(1)\ln\frac{4\pi\mu^{2}}{m^{2}_{1}}+\ln\frac{4\pi\mu^{2}}{m^{2}_{1}}+\psi(1)-\frac{1}{2}\right).$ (58)
@@
$\frac{1}{2}\left(\delta m^{2}+\frac{1}{2}\delta\lambda\phi^{2}+\delta\eta\phi+\delta\xi R\right)\times\int\frac{d^{d}k}{(2\pi)^{d}}\left(\frac{1}{k^{2}+m^{2}_{1}}+\frac{\left(\frac{1}{6}-\xi\right)R}{(k^{2}+m^{2}_{1})^{2}}\right),$ (59)
where $\delta m^{2}$, $\delta\lambda$ and $\delta\xi$ can be seen in 19. We next add the divergent terms coming from 57, 58, 59, in order to find out the divergence of the effective potential
@@
$\left.+\eta\lambda^{2}\phi^{3}\left(2\psi(2)-2\psi(1)-3\right)+\frac{\eta^{2}}{2}\left(\frac{1}{6}-\xi\right)R+\eta\lambda\phi\left(\frac{1}{6}-\xi\right)R+\frac{\lambda^{2}\phi^{2}}{2}\left(\frac{1}{6}-\xi\right)R\right).$ (60)
@@
$\delta\Lambda=-\frac{\mu^{-2\epsilon}\left(m^{4}\lambda+2m^{2}\eta^{2}\right)}{2(4\pi)^{4}\epsilon^{2}}-\frac{2\mu^{-\epsilon}m^{2}\eta^{2}\left(\psi(2)-\psi(1)-\frac{3}{2}\right)}{2(4\pi)^{4}\epsilon},$
$\delta\kappa=\frac{\mu^{-2\epsilon}\left(2m^{2}\lambda+\eta^{2}\right)}{2(4\pi)^{4}\epsilon^{2}}\left(\frac{1}{6}-\xi\right)-\frac{\eta^{2}\mu^{-\epsilon}}{4(4\pi)^{4}\epsilon}\left(\frac{1}{6}-\xi\right),$
$\delta m^{2}=\frac{\left(4m^{2}\lambda^{2}+\frac{7}{2}\lambda\eta^{2}\right)\mu^{-2\epsilon}}{(4\pi)^{4}\epsilon^{2}}+\frac{2m^{2}\lambda^{2}\mu^{-\epsilon}}{(4\pi)^{4}\epsilon}\left(\frac{3}{4}\psi(2)-\frac{1}{2}\psi(1)-\frac{3}{4}\right)+\frac{\eta^{2}\lambda\mu^{-\epsilon}}{(4\pi)^{4}\epsilon}\left(\frac{5}{2}\psi(2)-\frac{5}{2}\psi(1)-\frac{15}{4}\right),$
$\delta\xi=-\frac{\mu^{-2\epsilon}\lambda^{2}}{(4\pi)^{4}\epsilon^{2}}\left(\frac{1}{6}-\xi\right)+\frac{\lambda^{2}\mu^{-\epsilon}}{2(4\pi)^{4}\epsilon}\left(\frac{1}{6}-\xi\right),$
$\delta\tau=\frac{\left(4m^{2}\eta\lambda+\eta^{3}\right)\mu^{-2\epsilon}}{2(4\pi)^{4}\epsilon^{2}}+\frac{m^{2}\eta\lambda\mu^{-\epsilon}}{2(4\pi)^{4}\epsilon}\left(2\psi(2)-2\psi(1)-3\right)+\frac{\eta^{3}\mu^{-\epsilon}}{2(4\pi)^{4}\epsilon}\left(\psi(2)-\psi(1)-\frac{3}{2}\right),$
$\delta\gamma=-\frac{4\eta\lambda\mu^{-2\epsilon}}{2(4\pi)^{4}\epsilon^{2}}\left(\frac{1}{6}-\xi\right)-\frac{\eta\lambda\mu^{-\epsilon}}{2(4\pi)^{4}\epsilon}\left(\frac{1}{6}-\xi\right),$
$\delta\eta=\frac{9\mu^{-2\epsilon}\eta\lambda^{2}}{(4\pi)^{4}\epsilon^{2}}+\frac{3\mu^{-\epsilon}\eta\lambda^{2}}{(4\pi)^{4}\epsilon}\left(2\psi(2)-2\psi(1)-3\right),$
$\delta\lambda=\frac{9\lambda^{3}\mu^{-2\epsilon}}{(4\pi)^{4}\epsilon^{2}}+\frac{6\lambda^{3}\mu^{-\epsilon}}{(4\pi)^{4}\epsilon}\left(\psi(2)-\psi(1)-\frac{3}{2}\right).$ (61)
@@
$\left.+\ln\frac{4\pi\mu^{2}}{m^{2}_{1}}+\psi(1)-\frac{1}{2}\right).$ (62)
@@
$\left.+2\ln{\frac{4\pi\mu^{2}}{m^{2}_{1}}}\left(\psi(1)+\psi(2)\right)\right)+{\cal O}(\lambda^{2}),$ (63)
where the one loop effective potential is given by 21. However, we note that the one loop effective potential contains terms quadratic in the curvature as well, whereas in the two loop computation, we have retained terms only linear in the curvature. Hence we shall add with 63 the second order in curvature contribution coming from the ${\cal O}(\lambda)$ double bubble, reading
$V_{{\rm eff},{\cal O}(R^{2})}^{2-{\rm loop}}=-\frac{\lambda}{8(4\pi)^{4}}\left({2f_{1}}\left(\ln{\frac{m^{2}_{1}}{4\pi\mu^{2}}-\psi(2)}\right)+\left(\frac{1}{6}-\xi\right)^{2}R^{2}\left(4\psi(1)\ln{\frac{4\pi\mu^{2}}{m_{1}^{2}}}+\frac{3}{2}\left(\ln{\frac{4\pi\mu^{2}}{m_{1}^{2}}}\right)^{2}+\frac{\pi^{2}}{3}+2\psi^{2}(1)-\psi^{\prime}(1)\right)\right).$ (64)
We shall not go into the detail of the derivation for the above. The above contribution only required some gravitational counterterms for the sake of renormalisation. We have plotted 63 plus 64 in 16 for the de Sitter spacetime. Note that compared to the zero temperature one loop result, 44, there is no qualitatively new change. In particular for $\bar{\eta}=0$, the symmetry breaking feature remains intact.
@@
Figure 16: The two loop double bubble effective potential, 63, at vanishing temperature in the de Sitter spacetime for values $\bar{\eta}=0.447,\ \lambda=0.1,\ \xi=0.125,\ \Lambda=30{\rm GeV}^{2}$. Note that for the quartic self interaction, the symmetry breaking feature remains intact compared to that of one loop, 5. As earlier, our renormalisation scheme ensures that there are no $\phi$-independent terms in the effective potential.
@@ 6.1 Two loop at finite temperature
$\frac{1}{k^{2}+m^{2}}\rightarrow\frac{1}{k^{2}+m^{2}}+2\pi n_{B}({|k^{0}|})\delta(k^{2}-m^{2})=\frac{1}{k^{2}+m^{2}}+n_{1}(\beta,k)\qquad({\rm say}),$ (65)
and
$\frac{1}{\left(k^{2}+m^{2}\right)^{2}}\rightarrow\frac{1}{\left(k^{2}+m^{2}\right)^{2}}+n_{2}(\beta,k),$ (66)
@@
$\int\frac{d^{d}k}{(2\pi)^{d}}\frac{1}{k^{2}+m^{2}}\rightarrow\int\frac{d^{d}k}{(2\pi)^{d}}\frac{1}{k^{2}+m^{2}}+S_{1}(\beta)=-m^{2}\left(\frac{2\mu^{-\epsilon}}{\epsilon}+\psi(2)+\ln{\frac{4\pi\mu^{2}}{m^{2}}}\right)+S_{1}(\beta),$ (67)
@@
$\int\frac{d^{d}k}{(2\pi)^{d}}\frac{1}{(k^{2}+m^{2})^{2}}\rightarrow\int\frac{d^{d}k}{(2\pi)^{d}}\frac{1}{(k^{2}+m^{2})^{2}}+S_{2}(\beta)=\left(\frac{2\mu^{-\epsilon}}{\epsilon}+\psi(1)+\ln{\frac{4\pi\mu^{2}}{m^{2}}}\right)+S_{2}(\beta).$ (68)
@@
$\frac{\left(\eta+\lambda\phi\right)^{2}}{12}\left(\int\frac{d^{d}k_{1}}{(2\pi)^{d}}\frac{d^{d}k_{2}}{(2\pi)^{d}}\left(\frac{1}{k_{1}^{2}+m^{2}_{1}}+n_{1}(\beta,k_{1})\right)\left(\frac{1}{k_{2}^{2}+m^{2}_{1}}+n_{1}(\beta,k_{2})\right)\left(\frac{1}{(k_{1}+k_{2})^{2}+m^{2}_{1}}+n_{1}(\beta,k_{1}+k_{2})\right)\right.$
$\left.+\left(\frac{1}{6}-\xi\right)R\int\frac{d^{d}k_{1}}{(2\pi)^{d}}\frac{d^{d}k_{2}}{(2\pi)^{d}}\left(\frac{1}{k_{1}^{2}+m^{2}_{1}}+n_{1}(\beta,k_{1})\right)\left(\frac{1}{(k_{2}^{2}+m^{2}_{1})^{2}}+n_{2}(\beta,k_{2})\right)\left(\frac{1}{(k_{1}+k_{2})^{2}+m^{2}_{1}}+n_{1}(\beta,k_{1}+k_{2})\right)\right.$
@@
$+\frac{\left(\eta+\lambda\phi\right)^{2}}{12}\left(\frac{6\mu^{-\epsilon}}{\epsilon}S_{1}(\beta)+\left(\frac{1}{6}-\xi\right)R\frac{6\mu^{-\epsilon}}{\epsilon}S_{2}(\beta)\right)+\rm{Finite\,\,terms}.$ (70)
@@
$+\frac{\mu^{-\epsilon}}{2(4\pi)^{2}\epsilon}\left(-m^{2}\lambda-\eta^{2}-\frac{3}{2}\lambda^{2}\phi^{2}-3\eta\lambda+\left(\frac{1}{6}-\xi\right)\lambda R\right)\left(S_{1}(\beta)+\left(\frac{1}{6}-\xi\right)RS_{2}(\beta)\right).$ (71)
We now collect and add all the divergent terms from 69, 70 and 71. We obtain after some calculations,
@@
$\left.+\left(\frac{1}{6}-\xi\right)R\int\frac{d^{d}k_{1}}{(2\pi)^{d}}\frac{d^{d}k_{2}}{(2\pi)^{d}}\frac{1}{k_{1}^{2}+m^{2}_{1}}\frac{1}{k_{2}^{2}+m^{2}_{1}}\frac{1}{((k_{1}+k_{2})^{2}+m^{2}_{1})^{2}}\right),$ (72)
@@
$\left.-\frac{15m_{1}^{6}\beta^{4}\zeta(7)}{3072\pi^{8}}+\frac{15m_{1}^{7}\beta^{5}\zeta(7)}{24576\pi^{9}}-\frac{15m_{1}^{7}\beta^{5}\zeta(7)}{393216\pi^{9}}+\frac{3m_{1}^{8}\beta^{6}\zeta(3)\zeta(5)}{524288\pi^{10}}+\frac{m_{1}^{8}\beta^{6}\zeta(3)\zeta(5)}{262144\pi^{10}}\right)+{\cal O}(R^{2}),$ (73)
where $V^{(2)}_{\rm{db}}$ is the corresponding zero temperature part and the cross terms of zero and finite temperature. Inclusion of ${\cal O}(R^{2})$ terms to the above result is easy, but we shall not go into such computation here.